A cubic sample of a new kind of artificial tissue is subject to an increase in pressure of 160 kPa which results in a reduction in the side length of the cube of 6.5%. That is, the side length of the cube has reduced from L0 to 0.935L0.
what is the bulk modulus
1
Expert's answer
2020-04-17T10:48:09-0400
"K=\frac{\sigma}{\epsilon_V}=\frac{\sigma}{3\epsilon_L}"

"K=\frac{160}{3\cdot 0.065}=820\ kPA=0.82\ MPa"